OpenStudy (anonymous):

What is the solution to the equation? 32 - 6x + 7x = -15 + 28 X = ????

OpenStudy (anonymous):

@callmechris99

OpenStudy (anonymous):

X= -19 Post this in the math section next time!

OpenStudy (anonymous):

32 - 6x + 7x = -15 + 28 or, 32+x=13 or, x= 32-13 or, x=19 . !9 is the right answer . You can give me a medal . Thanx :)